# UQLM: A Python Package for Uncertainty Quantification in Large Language Models

### v1

@@ Abstract
Hallucinations, defined as instances where Large Language Models (LLMs) generate false or misleading content, pose a significant challenge that impacts the safety and trust of downstream applications. We introduce uqlm, a Python package for LLM hallucination detection using state-of-the-art uncertainty quantification (UQ) techniques. This toolkit offers a suite of UQ-based scorers that compute response-level confidence scores ranging from 0 to 1. This library provides an off-the-shelf solution for UQ-based hallucination detection that can be easily integrated to enhance the reliability of LLM outputs. [^1] Note: This paper is not intended to reflect the specific work, practices, or opinions of the author-affiliated company. Any resemblance to actual practices is coincidental. All opinions are the authors’ own.
@@ 1 Introduction
Existing solutions to this problem include source-comparison methods, internet-based grounding, and uncertainty quantification (UQ) methods. Toolkits that offer source-comparison scorers, such as Ragas (Es et al., 2023), Phoenix (Arize AI, 2025), DeepEval (Ip and Vongthongsri, 2025), and others (Hu et al., 2024; UpTrain AI Team, 2024; Zha et al., 2023; Asai et al., 2023) evaluate the consistency between generated content and input prompts. However, these methods can mistakenly validate responses that merely mimic prompt phrasing without ensuring factual accuracy. Toolkits that leverage Internet searches for fact-checking, such as FacTool (Chern et al., 2023), introduce delays and risk incorporating erroneous online information, failing to address the inherent uncertainty in model outputs. Lastly, although numerous UQ techniques have been proposed in the literature, their adoption in user-friendly, comprehensive toolkits remains limited. For example, while SelfCheckGPT (Manakul et al., 2023) incorporates some UQ scorers, its set of techniques is narrow and does not integrate generation with evaluation, thus creating barriers for practitioners outside specialized AI research environments. LangKit (WhyLabs, 2025) and NeMo Guardrails (Rebedea et al., 2023) also offer UQ scorers but are similarly narrow in scope. Lastly, while LM-Polygraph (Fadeeva et al., 2023) offers a robust collection of UQ-based approaches for LLMs, its documentation is written for a research audience and may lack ease of use for non-specialized practitioners.
We aim to bridge these gaps by introducing a comprehensive open-source Python package, uqlm, that democratizes advanced research in LLM uncertainty quantification. UQLM (Uncertainty Quantification for Language Models) implements a diverse array of uncertainty estimation techniques to compute generation-time, response-level confidence scores and uniquely integrates generation and evaluation processes. This integrated approach allows users to generate and assess content simultaneously, without the need for ground-truth data or external knowledge sources, and with minimal engineering effort. This democratization of access empowers smaller teams, researchers, and developers to incorporate robust hallucination detection into their applications, contributing to the development of safer and more reliable AI systems.
@@ 2 Usage
The uqlm library, available at https://github.com/cvs-health/uqlm, provides a collection of UQ-based scorers spanning four categories: black-box UQ, white-box UQ, LLM-as-a-Judge, and ensembles. [^2] For an detailed overview of available scorers and associated experiment results, we refer to the reader to this project’s companion paper, Bouchard and Chauhan (2025). The corresponding classes for these techniques are instantiated by passing an LLM object to the constructor. [^3] For the current version of uqlm, v0.1.8, a LangChain BaseChatModel is required. Note that an LLM is not required if users provide pre-generated responses and implement the score method. Each of these classes contains a generate_and_score method, which generates LLM responses to a user provided list of prompts and computes response-level confidence scores, which range from 0 to 1.
@@ 2.1 Black-Box Uncertainty Quantification
Black-box uncertainty quantification exploits the stochastic nature of LLMs and measures the consistency of multiple responses to the same prompt. These consistency measurements can be conducted with various approaches, including semantic entropy (Farquhar et al., 2024), non-contradiction probability (Chen and Mueller, 2024; Lin et al., 2024), BERTScore (Manakul et al., 2023; Zha et al., 2023), BLEURT (Sellam et al., 2020), exact match rate (Cole et al., 2023), and cosine similarity (Shorinwa et al., 2024). Black-box UQ scorers are compatible with any LLM, but increase latency and generation costs. The corresponding class for this collection of scorers is BlackBoxUQ.
Figure 1: Illustration of a Black-Box Scorer Workflow
To implement BlackBoxUQ.generate_and_score, users provide a list of prompts. For each prompt, an original response, along with additional candidate responses, are generated by the user-provided LLM, and consistency scores are computed using the specified scorers (see Figure 1). [^4] Note that Figure 1 depicts the approach for all black-box UQ scorers except semantic entropy, which does not designate an ‘original response’. If users set use_best=True, the uncertainty-minimized response is selected. [^5] Uncertainty-minimized response selection is based on semantic entropy (Farquhar et al., 2024). Below is a minimal example illustrating usage of BlackBoxUQ.
@@ 2.2 White-Box Uncertainty Quantification
White-box uncertainty quantification leverages token probabilities to compute uncertainty, as depicted in Figure 2. These approaches have the advantage of using the token probabilities associated with the generated response, meaning they do not add any latency or generation cost. However, because token probabilities are not accessible from all APIs, white-box scorers may not be compatible with all LLM applications. This collection of scorers can be implemented with the WhiteBoxUQ class, which offers two scorers: minimum token probability (Manakul et al., 2023) and length-normalized token probability (Malinin and Gales, 2021). Below is a minimal example of WhiteBoxUQ usage.
@@
Figure 2: Illustration of a White-Box Scorer Workflow
@@ 2.3 LLM-as-a-Judge
LLM-as-a-Judge uses an LLM to evaluate the correctness of a response to a particular question. To achieve this, a question-response concatenation is passed to one or more LLMs along with instructions to score the response’s correctness using the LLMPanel class (see Figure 3). In the constructor, users pass a list of LLM objects to the judges argument and specify one of four scoring templates for each judge with the scoring_templates argument. These four scoring templates are as follows: binary ($\{\text{incorrect},\text{correct}\}{ incorrect , correct }$ as $\{0,1\}{ 0 , 1 }$), ternary ($\{\text{incorrect},\text{uncertain},\text{correct}\}{ incorrect , uncertain , correct }$ as $\{0,0.5,1\}{ 0 , 0.5 , 1 }$), continuous (any value between 0 and 1), and a 5-point Likert scale (0,0.25,,100.2510,0.25,...,10 , 0.25 , … , 1). Implementing the generate_and_score method returns the score from each judge and aggregations of these scores, including minimum, maximum, average, and median. See below for a minimal example.
Figure 3: Illustration of LLM-as-a-Judge Workflow
@@ Acknowledgments
We wish to thank Piero Ferrante, Blake Aber, Matthew Churgin, Erik Widman, Robert Enzmann, and Xue (Crystal) Gu for their helpful suggestions.
Appendix A Ensemble Tuning.
In order to tune the ensemble weights prior to using the generate_and_score method, users must provide a list of prompts and corresponding ideal responses to serve as an ‘answer key’. The LLM’s responses to the prompts are graded with a grader function that compares against the provided ideal responses. If a grader function is not provided by the user, the default grader function that leverages vectara/hallucination_evaluation_model is used.

### v2

@@ Abstract
Hallucinations, defined as instances where Large Language Models (LLMs) generate false or misleading content, pose a significant challenge that impacts the safety and trust of downstream applications. We introduce uqlm, a Python package for LLM hallucination detection using state-of-the-art uncertainty quantification (UQ) techniques. This toolkit offers a suite of UQ-based scorers that compute response-level confidence scores ranging from 0 to 1. This library provides an off-the-shelf solution for UQ-based hallucination detection that can be easily integrated to enhance the reliability of LLM outputs.
@@ 1 Introduction
Existing solutions to this problem include source-comparison methods, internet-based grounding, and uncertainty quantification (UQ) methods. Toolkits that offer source-comparison scorers, such as Ragas (Es et al., 2023), Phoenix (Arize AI, 2025), DeepEval (Ip and Vongthongsri, 2025), and others (Hu et al., 2024; UpTrain AI Team, 2024; Zha et al., 2023; Asai et al., 2023) evaluate the consistency between generated content and input prompts. However, these methods can mistakenly validate responses that merely mimic prompt phrasing without ensuring factual accuracy. Toolkits that leverage Internet searches for fact-checking, such as FacTool (Chern et al., 2023), introduce delays and risk incorporating erroneous online information, failing to address the inherent uncertainty in model outputs. Lastly, although numerous UQ techniques have been proposed in the literature, their adoption in user-friendly, comprehensive toolkits remains limited. For example, while SelfCheckGPT (Manakul et al., 2023) incorporates some UQ scorers, its set of techniques is narrow and does not integrate generation with evaluation, thus creating barriers for practitioners outside specialized AI research environments. LangKit (WhyLabs, 2025) and NeMo Guardrails (Rebedea et al., 2023) also offer UQ scorers but are similarly narrow in scope. LM-Polygraph (Fadeeva et al., 2023) provides a valuable collection of UQ-based approaches for LLMs and is built on the Hugging Face ecosystem, representing an important step toward more accessible UQ tools, though opportunities remain for enhancing accessibility for non-specialized practitioners.
We aim to bridge these gaps by introducing a comprehensive open-source Python package, uqlm, that democratizes advanced research in LLM uncertainty quantification. UQLM (Uncertainty Quantification for Language Models) implements a diverse array of uncertainty estimation techniques to compute generation-time, response-level confidence scores and uniquely integrates generation and evaluation processes. This integrated approach allows users to generate and assess content simultaneously, without the need for ground-truth data or external knowledge sources, and with minimal engineering effort. This democratization of access empowers smaller teams, researchers, and developers to incorporate robust hallucination detection into their applications for safer and more reliable AI systems.
@@ 2 Usage
The uqlm library, available at https://github.com/cvs-health/uqlm, provides a collection of UQ-based scorers spanning four categories: black-box UQ, white-box UQ, LLM-as-a-Judge, and ensembles. [^1] For a detailed overview of available scorers and associated experiment results, we refer the reader to this project’s companion paper, Bouchard and Chauhan (2025). The corresponding classes for these techniques are instantiated by passing a LangChain BaseChatModel (LLM) instance to the constructor. [^2] UQLM is compatible with all LangChain Chat Models. More details are provided in Appendix B. Each of these classes contains a generate_and_score method, which generates LLM responses to a user provided list of prompts and computes response-level confidence scores, which range from 0 to 1.
@@ 2.1 Black-Box Uncertainty Quantification
Black-box uncertainty quantification exploits the stochastic nature of LLMs and measures the consistency of multiple responses to the same prompt. These consistency measurements can be conducted with various approaches, including discrete semantic entropy (Farquhar et al., 2024), number of semantic sets (Lin et al., 2024), non-contradiction probability (Chen and Mueller, 2024), entailment probability (Lin et al., 2024), BERTScore (Manakul et al., 2023), exact match rate (Cole et al., 2023), and cosine similarity (Shorinwa et al., 2024). Black-box UQ scorers are compatible with any LLM, but increase latency and generation costs. The BlackBoxUQ class uses the user-provided LLM to generate, for each prompt, an original response and additional candidate responses, then computes consistency scores using the specified scorers (see Figure 1).  [^3] Users may skip generation by calling the score method with pre-generated responses. The workflow depicted in Figure 1 does not apply to semantic entropy, which does not designate an ‘original response’. See below for a minimal BlackBoxUQ example.444use_best=True selects the mode from the highest-probability semantic cluster (Farquhar et al., 2024).
@@ 2.2 White-Box Uncertainty Quantification
White-box uncertainty quantification leverages token probabilities to compute uncertainty, as in Figure 2. UQLM’s WhiteBoxUQ class includes several categories of methods. Single-generation methods, which offer the advantage of no additional latency or generation costs, include minimum token probability (Manakul et al., 2023), length-normalized token probability (Malinin and Gales, 2021), sequence probability (Vashurin et al., 2025a), likelihood margin (Farr et al., 2024), mean top-K token entropy (Scalena et al., 2025), and maximum top-K token entropy (Scalena et al., 2025). Sampling-based methods include semantic entropy (Farquhar et al., 2024), semantic density (Qiu and Miikkulainen, 2024), Monte Carlo predictive entropy (Kuhn et al., 2023), and CoCoA (Vashurin et al., 2025b). Additionally, the class implements the $P(\text{True})$ method (Kadavath et al., 2022), which requires one extra generation per response. Note that WhiteBox UQ will only work with model APIs that expose token probabilities. See below for example WhiteBoxUQ usage.
@@
Figure 2: Illustration of a Single-Generation White-Box Scorer Workflow
@@ 2.3 LLM-as-a-Judge
LLM-as-a-Judge uses an LLM to evaluate the correctness of a response to a particular question. To achieve this, a question-response concatenation is passed to one or more LLMs along with instructions to score the response’s correctness using the LLMPanel class (see Figure 3). In the constructor, users pass a list of LLM objects to the judges argument and specify one of four scoring templates for each judge with the scoring_templates argument. These four scoring templates are as follows: binary ($\{\text{incorrect},\text{correct}\}$ as $\{0,1\}$), ternary ($\{\text{incorrect},\text{uncertain},\text{correct}\}$ as $\{0,0.5,1\}$), continuous (any value between 0 and 1), and a 5-point Likert scale (0,0.25,,10,0.25,...,1). Instructions are customized using the additional_context argument. Implementing the generate_and_score method (minimal example below) returns the score from each judge along with aggregations such as average, minimum, and median.
@@ Acknowledgments
We wish to thank Piero Ferrante, Blake Aber, Matthew Churgin, Erik Widman, Robert Enzmann, Xue (Crystal) Gu and Huiwen Hu for their helpful suggestions as well as Dimitris Tsapetis, Vipin Gyanchandani, Drew Ross, Jinesh Mehta, Namrata Walanj, and Joshua Mabry for their contributions to UQLM.
Appendix A Ensemble Tuning
In order to tune the ensemble weights prior to using the generate_and_score method, users must provide a list of prompts and corresponding ideal responses to serve as an ‘answer key’. The LLM’s responses to the prompts are graded with a grader function that compares against the provided ideal responses. If a grader function is not provided by the user, the original LLM is used as an LLM-based grader to evaluate the correctness of the responses relative to the answer key.
@@ Appendix A Ensemble Tuning
Appendix B Compatible Models
UQLM integrates with the LangChain Chat Model interface, making it compatible with a large and evolving ecosystem of providers and models; an exhaustive list is infeasible. In practice, users access models through LangChain wrappers for commercial API providers such as OpenAI, Anthropic, Google, Cohere, and AWS Bedrock, as well as for locally deployed open-source models like Llama 3, Mistral, and Qwen using tools such as Ollama. For an up-to-date catalog and provider-specific options, we refer readers to the LangChain documentation for supported chat models. [^5] https://python.langchain.com/docs/integrations/chat/ Note that white-box uncertainty methods require access to token log probabilities, which some wrappers do not expose. LangChain documents how to request these log probabilities and how to verify whether a specific wrapper returns them. [^6] https://python.langchain.com/docs/how_to/logprobs/ Table 1 provides detailed compatibility information for each scorer family.
Scorer family Needs logprobs Compatibility
Black-Box UQ No Compatible with all chat models
White-Box UQ Yes Chat models that expose logprobs
LLM-as-a-Judge No Compatible with all chat models
UQ Ensemble Depends Depends on included scorers
Table 1: UQLM method families and interface requirements.